OpenStudy (anonymous):

Matilda practices guitar for a total of 70 minutes every day. She spends 10 minutes more practicing the electric guitar than the acoustic guitar. Write a pair of linear equations to show the relationship between the number of minutes Matilda practices the acoustic guitar every day (x) and the number of minutes she practices the electric guitar every day (y). x + y = 70 x + 10 = y How much time does Matilda spend on practicing the acoustic guitar every day?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Ive been needing help with this for hours..

OpenStudy (anonymous):

i already have the linear equations i just need help with the other one

OpenStudy (texaschic101):

x + 10 = y so sub in x + 10 for y in the other equation x + y = 70 x + x + 10 = 70 2x + 10 = 70 2x = 60 x = 30 x + 10 = y 30 + 10 = y 40 = y acoustic (x) = 30 electric (y) = 40

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Thank you! and yes there is one last part Is it possible for Matilda to have spent 50 minutes practicing the electric guitar every day? Explain your reasoning.

OpenStudy (texaschic101):

no....because then she would be spending 50 on electric and 20 on acoustic....a difference of 30 minutes. Or she would have to allow for more time in practices

OpenStudy (texaschic101):

wait...I gues it would be possible. She could spend 50 on electric and 20 on acoustic. Is it possible ? I suppose it is if you are allowed to change things.

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

Its possible if she can change the time limit on acoustic. If she can't, then it is not possible

OpenStudy (texaschic101):

happy to help :)